# ADF22+: a declining faint end in the far-infrared luminosity function in the SSA22 protocluster at z=3.09

### v1

@@
Protoclusters represent the densest regions of cosmic large-scale structure in the early universe and are the environment where present-day massive elliptical galaxies are assembled. Millimeter continuum emission offers a powerful probe of obscured star formation at high redshifts across various environments. In this paper, we present deep ALMA 1.17 mm mosaic of the central 8 arcmin2 ( comoving Mpc2) region in the SSA22 protocluster at to study the faint dusty star-forming galaxy (DSFG) population. The continuum map achieves an RMS noise level of Jy beam-1 at spatial resolution, the depth of previous observation of this field. We detected 53 sources with a signal-to-noise ratio above 4.2, doubling the number of detections. Utilizing optical to mid-infrared ancillary data, we search for spectroscopic redshift and identify 18 of 53 as cluster members. For sources with more than two photometric data points in the near-infrared, stellar mass () and star formation rate (SFR) from spectral energy distribution fitting are presented. The 1.17 mm number count shows excess at flux density mJy but are consistent with blank field in fainter flux bins. The monochromatic far-infrared luminosity function of the SSA22 protocluster core region suggests a lack of faint DSFGs. All SSA22 protocluster member galaxies detected at 1.17 mm have SFR within the -SFR relation of general star-forming galaxies. Our results suggest that an early overdense environment like SSA22 protocluster predominantly enhances the formation of massive early-type galaxies in present-day galaxy clusters, but that the star formation in individual member galaxies is likely driven by gas supply along the cosmic web and occurs in a secular way.
@@
Overdense regions in the high-redshift universe are thought to be the progenitor of present-day galaxy clusters and are referred to as protoclusters (e.g., Overzier 2016). In CDM cosmology, structure formation occurs in a hierarchical manner. Peaks in the initial density fluctuation collapse into dark matter halos (Gunn & Gott 1972) and define the filaments and sheets in the large-scale matter distribution (the “cosmic web”, Bond et al. 1996). The most massive halos form from mergers of smaller ones at the intersecting nodes of cosmic filaments (e.g., Boylan-Kolchin et al. 2009), where large amounts of infalling materials along the cosmic web are thought to fuel star formation and give rise to protocluters of galaxies in the early universe (Umehata et al. 2019). Observations of protoclusters suggest that galaxy evolution is accelerated in them compared to the field, signposted by elevated levels of star formation (e.g., Stevens et al. 2003; Dannerbauer et al. 2014; Casey et al. 2015; Miller et al. 2018), higher active galactic nuclei fractions (AGN, e.g., Lehmer et al. 2009a, b, 2013; Tozzi et al. 2022), and eventually, early emergence of massive quiescent galaxies (e.g., Kubo et al. 2021; McConachie et al. 2022; Ito et al. 2023). The rapid growth and quenching of massive galaxies in protoclusters can explain old red giant elliptical galaxies in the center of present-day clusters (e.g., Nelan et al. 2005).
@@
Most of the star formation in massive galaxies at high redshift is obscured by dust ( for log, e.g., Whitaker et al. 2017; Dudzevičiūtė et al. 2020), thus one of the most efficient ways to probe their stellar mass build-up is to observe the rest-frame far-infrared emission which originates from dust heating by young massive stars. There are numerous studies targeting high-redshift overdensities using single-dish submillimeter/millimeter (submm/mm) telescopes equipped with bolometer cameras (e.g., Ivison et al. 2000; Geach et al. 2005; Tamura et al. 2009; Chapman et al. 2009; Umehata et al. 2014; Dannerbauer et al. 2014; Clements et al. 2016; Arrigoni Battaia et al. 2018; Smail 2024; Wang et al. 2024). These observations have consistently found abundant bright submm/mm sources indicating intense star formation enshrouded by dust in protoclusters. While single-dish surveys provide a wide field-of-view (FOV) which is needed to map protocluster fields, the coarse angular resolution limits the achievable sensitivity due to confusion noise and complicates the identification of counterparts. Interferometers offer angular resolution high enough to overcome the confusion limit of single-dish telescopes, but the small FOV necessitates mosaicking, so interferometric protocluster surveys with a significant area are scarce. Existing interferometric surveys of protoclusters (e.g., Umehata et al. 2017; Oteo et al. 2018) are often too shallow to compare with deep blank field mosaics (e.g., González-López et al. 2020; Fujimoto et al. 2024). As such, the effect of protocluster environment on the abundance of fainter and more representative populations of dusty star-forming galaxies remains poorly explored. Investigating the dust content of other populations in protocluster such as Ly blobs (e.g., Umehata et al. 2021) and Lyman break galaxies also requires substantially deeper data.
@@
The Special Selected Area 22 (SSA22) protocluster at was first discovered as a peak in redshift distribution of Lyman break galaxies (Steidel et al. 1998), and then associated with a Mpc scale structure identified by overdensity of Ly emitters (Hayashino et al. 2004; Yamada et al. 2012) distributed along at least three filaments (Matsuda et al. 2005). The large-scale filaments intersect at the core region of the SSA22 protocluster, where numerous massive galaxies (Kubo et al. 2013, 2015, 2016) and AGNs (Lehmer et al. 2009a, b) are seen. Since its discovery, it has been extensively studied from X-ray to radio wavelengths (e.g., Chapman et al. 2004a; Hayashino et al. 2004; Lehmer et al. 2009b; Webb et al. 2009; Kato et al. 2016; Ao et al. 2017). Single-dish submm/mm surveys of the SSA22 field (Geach et al. 2005; Tamura et al. 2009; Umehata et al. 2014) have found substantial star formation activities obscured by dust. Umehata et al. (2017) mapped the central 7 arcmin2 ( Mpc2) area at 1.1 mm using ALMA and revealed an increased number density of bright dusty star-forming galaxies (DSFGs) in the core region. The distribution of member DSFGs and AGNs overlaps with diffuse gas filaments seen in Ly emission (Umehata et al. 2019), suggesting galaxy growth fueled by gas supply from the cosmic web. However, the abundance of fainter DSFGs was not clear because of the small number of detection at an RMS sensitivity of 60 Jy beam-1 (equivalent to a SFR limit of yr-1 at ).
@@ 2.1 ALMA band 6 continuum mosaic
The data were processed by ALMA pipeline in the CASA (CASA Team et al. 2022) package. Imaging is performed using tclean task with Briggs weighting (robust=0.5), and auto-multithresh masking. We choose the mtmfs deconvolver with scale=[0] because the continuum sources are expected to be at most marginally resolved. The resulting 256.98 GHz (1.17 mm) continuum image is shown in Fig. 1. The synthesized beam has FWHM and position angle (PA)=. The RMS noise level is 25 Jy beam-1 before primary beam correction, of the depth of Jy beam-1 map from Umehata et al. (2017).
@@ 3.1 Source detection
In Fig 2a, we show the false detection rate as a function of peak S/N. Based on these results, we filter the initial source candidates by requiring peak S/N and create the final source catalogue (Appendix A, see also Fig. 1) from 53 sources, corresponding to a spurious rate of . For each detection, we search for a counterpart in Umehata et al. (2017) main and supplementary catalogues (A1–A32) using a matching radius of beam FWHM (, equivalent to of Gaussian). If no match is found, a new ID starting from A33 is assigned in the order of descending flux. All 18 sources in the main catalogue of Umehata et al. (2017) are detected in the new deeper mosaic, but only 3 out of 14 in the supplementary catalogue with S/N–5.0 (or 2 of 8 sources with S/N) are recovered.
Source position and flux are measured using the CASA imfit task. The detection completeness is estimated by injecting point sources to random blank positions in the image before primary beam correction, and then running the same source detection script. The artificial source is considered as recovered if it is detected at peak S/N within from the injected position. The completeness as a function of source flux is shown in Figure 2b. From the results, we also calculate the ratio between measured and true fluxes (Figure 2c). For sources with peak S/N, we divide their 1.17 mm fluxes from imfit task by the median measured-to-true flux ratio at the corresponding S/N to correct for flux boosting.
@@ 3.2 Counterpart identification and source properties
To search for counterparts at optical and near-infrared wavelengths, we first align the images to the ALMA coordinate frame and subtract the background. A gallery of image stamps at ALMA source position is provided in Appendix C. The ground-based images have a homogenized point spread function (PSF) FWHM of . Photometry on these images is performed in diameter circular aperture centered on the ALMA position and corrected for aperture loss using the point spread function model in each band. Flux errors are estimated as the standard deviations of fluxes from 1000 randomly placed apertures at blank positions. Fluxes in Spitzer IRAC bands are measured from point source decomposition of the image (Hsieh et al. 2012) in the same aperture as optical images, which allows the source centroid to wander by up to . In Table 1, sources with S/N in any of the , IRAC1 and IRAC2 bands (median depth = 0.3 Jy, 0.3 Jy and 0.4 Jy) are marked as NIR detected. Three sources (A33, A58 and A64) have a nearby NIR source located away from ALMA position and no photometry. They are marked as potential lensed sources or no counterpart (L/N) in our catalogue.
@@
The effective area as a function of flux is defined as the area in the mosaic where a source can be detected with S/N as shown in Figure 2b. We evaluate the number count from 0.06 to 6.3 mJy in six bins placed evenly in logarithmic space. Field-only number counts are also calculated by assigning a weight of zero to confirmed SSA22 members with spectroscopic redshift . To estimate uncertainty, we resample the flux column of the source table (Appendix A) by 1000 times, assuming flux error follows Gaussian distribution, and add the resulting standard deviation of number count in quadrature to the Poisson error as final uncertainty. The results are listed in Table 1 and plotted in Figure 4, together with a collection of literature results of blank fields (Karim et al. 2013; Ono et al. 2014; Carniani et al. 2015; Hatsukade et al. 2016, 2018; Stach et al. 2018; González-López et al. 2020; Simpson et al. 2020; Béthermin et al. 2020; Gómez-Guijarro et al. 2022; Chen et al. 2023; Fujimoto et al. 2024) and protocluster field (Zhang et al. 2024). The literature fluxes are scaled to 1.17 mm by assuming a modified black body with dust temperature K, emissivity , and median redshift of the DSFG population (e.g., Gómez-Guijarro et al. 2022; Dudzevičiūtė et al. 2020). The resulting number count shows consistency with field average for flux density mJy but obvious excess at brighter fluxes.
With a similar Monte Carlo method, we also compute the rest-frame 250 m luminosity function for the core region of the SSA22 protocluster (, Matsuda et al. 2005). Assuming a modified black body with dust temperature K, emissivity , and redshift , the 1.17 mm (rest-frame 285 m) luminosity is divided by 0.72 to convert to that at rest-frame 250 m. Protocluster membership is based on spectroscopic redshifts or the probability of photometric redshifts (see the next section) falling into the range of the core region. We resample the catalogue by 40,000 times and find the probability that one or more photometric redshift sources are protocluster members to be . However, we note that membership based on photometric redshift can be unreliable: the two sources (A34 and A49) with photometric redshift closest to 3.09 lie in deep CO(3-2) footprint but are undetected even their continuum fluxes are comparable to A42 at . As such, we mainly consider the results based on spectroscopic membership. The monochromatic luminosity function is listed in Table 2 and plotted in Fig 4.
@@ 4.2 The rest-frame FIR luminosity function
In Fig. 4, we can see that the rest-frame 250 m luminosity function of the SSA22 protocluster core region is dex above that in coeval blank field (e.g., Lapi et al. 2011). Such orders of magnitude higher obscured star formation rate density than field have also been observed in other high-redshift protoclusters (e.g., Oteo et al. 2018; Lacaille et al. 2019; Hill et al. 2020; Zhang et al. 2024). In line with what has been seen in number count, the luminosity function shows evidence of a declining (or at least flat) faint end, suggesting a dearth of faint DSFGs in the SSA22 protocluster core region: if the functional form is the same as blank fields, then by scaling the field luminosity function to match the two brightest flux bins of SSA22, we will expect much more sources at S mJy than what has been observed. The overall shape of the rest-frame 250 m luminosity function of the SSA22 protocluster core region is reminiscent of the Gaussian-shape bright end of the optical luminosity function of low-redshift galaxy clusters (e.g., Lobo et al. 1997; Wilson et al. 1997; Bernardi et al. 2003). This points to an evolutionary link between protocluster and local galaxy clusters and supports the idea that the massive early-type galaxies in local clusters have formed the bulk of their stellar mass during star formation events enshrouded by dust at (e.g., Stach et al. 2017; Cooke et al. 2019; Smail 2024). Finally, it is notable that recent high-resolution JWST and ALMA observations of DSFGs in the SSA22 protocluster core have revealed evidence of growing bulge components behind dust obscuration (Umehata et al. 2025), which again suggests that we are witnessing the early formation phase of a massive cluster.
@@ 4.3 Stellar mass and SFR of the SSA22 protocluster member galaxies
In Fig. 5, the SSA22 protocluster member galaxies have SFR overlapping with the SFMS, but the distributions of and SFR are different from field DSFGs. The differences are at significance levels from a two-sided KS test except for the distributions of SSA22 protocluster member and GOODS-ALMA DSFGs, which differs at . Overall, SSA22 protocluster galaxies have lower SFR and smaller , and we note that it is unlikely that the differences are driven by the choice of SED fitting codes because their results usually agree within error (Pacifici et al. 2023). Instead, it can be driven by the fainter fluxes of our sample: the median 1.17 mm fluxes are 0.8, 1.3 and 2.0 mJy for SSA22 protocluster member, GOODS-ALMA and AS2UDS+AS2COSMOS DSFGs, respectively. The deeper flux limit of our survey can pick up lower SFR and less massive systems, and the median of each sample (, , ) roughly scales as the median 1.17 mm fluxes. Since the submm/mm fluxes reflects a combination of dust mass and SFR (), this is the expected behavior if we further assume and consider the fact that none of the SSA22 protocluster member galaxies have SFR above SFMS like the field DSFGs. As is a proxy to total gas mass (e.g., Scoville et al. 2016; Hodge & da Cunha 2020), the above results suggest the current star formation in SSA22 protocluster member galaxies occurs in a self-regulated way powered by a large amount of fresh gas supply at the cosmic web node (Umehata et al. 2019), rather than short-lived starbursts induced by major merger. In addition, as will be presented in a future paper, [C ii] observations reveal that majority of bright SSA22 protocluster member DSFGs are regular rotating disks with wide-spread star formation (Umehata et al. 2025) instead of late-stage major merger, and the brightest one, A1, is a massive spiral galaxy (Umehata et al. 2024). These observations further support our argument about secular star formation. A similar conclusion has also been drawn for the Spiderweb protocluster (Pérez-Martínez et al. 2024).

### v2

@@
Protoclusters represent the densest regions of cosmic large-scale structure in the early universe and are the environment where present-day massive elliptical galaxies are assembled. Millimeter continuum emission offers a powerful probe of obscured star formation at high redshifts across various environments. In this paper, we present a deep ALMA 1.17 mm mosaic of the central 8 arcmin2 ( comoving Mpc2) region in the SSA22 protocluster at to study the faint dusty star-forming galaxy (DSFG) population. The continuum map achieves an RMS noise level of Jy beam-1 at spatial resolution, the depth of previous observation of this field. We detected 53 sources with a signal-to-noise ratio above 4.2, doubling the number of detections. Utilizing optical to mid-infrared ancillary data, we search for spectroscopic redshift and identify 18 of 53 as cluster members. For sources with more than two photometric data points in the near-infrared, stellar mass () and star formation rate (SFR) from spectral energy distribution fitting are presented. The 1.17 mm number count shows excess at flux density mJy but are consistent with blank field in fainter flux bins. The monochromatic far-infrared luminosity function of the SSA22 protocluster core region suggests a lack of faint DSFGs. All SSA22 protocluster member galaxies detected at 1.17 mm have SFR within the -SFR relation of general star-forming galaxies. Our results suggest that an early overdense environment like SSA22 protocluster predominantly enhances the formation of massive early-type galaxies in present-day galaxy clusters, but that the star formation in individual member galaxies is likely driven by gas supply along the cosmic web and occurs in a secular way.
@@
Overdense regions in the high-redshift universe are thought to be the progenitors of present-day galaxy clusters and are referred to as protoclusters (e.g., Overzier 2016). In CDM cosmology, structure formation occurs in a hierarchical manner. Peaks in the initial density fluctuation collapse into dark matter halos (Gunn & Gott 1972) and define the filaments and sheets in the large-scale matter distribution (the “cosmic web”, Bond et al. 1996). The most massive halos form from mergers of smaller ones at the intersecting nodes of cosmic filaments (e.g., Boylan-Kolchin et al. 2009), where large amounts of infalling materials along the cosmic web are thought to fuel star formation and give rise to protocluters of galaxies in the early universe (Umehata et al. 2019). Observations of protoclusters suggest that galaxy evolution is accelerated in them compared to the field, signposted by elevated levels of star formation (e.g., Stevens et al. 2003; Dannerbauer et al. 2014; Casey et al. 2015; Miller et al. 2018), higher active galactic nuclei fractions (AGN, e.g., Lehmer et al. 2009a, b, 2013; Tozzi et al. 2022), and eventually, early emergence of massive quiescent galaxies (e.g., Kubo et al. 2021; McConachie et al. 2022; Ito et al. 2023). The rapid growth and quenching of massive galaxies in protoclusters can explain old red giant elliptical galaxies in the center of present-day clusters (e.g., Nelan et al. 2005).
@@
Most of the star formation in massive galaxies at high redshift is obscured by dust ( for log, e.g., Whitaker et al. 2017; Dudzevičiūtė et al. 2020), thus one of the most efficient ways to probe their stellar mass build-up is to observe the rest-frame far-infrared emission which originates from dust heating by young massive stars. There are numerous studies targeting high-redshift overdensities using single-dish submillimeter/millimeter (submm/mm) telescopes equipped with bolometer cameras (e.g., Ivison et al. 2000; Geach et al. 2005; Tamura et al. 2009; Chapman et al. 2009; Umehata et al. 2014; Dannerbauer et al. 2014; Clements et al. 2016; Arrigoni Battaia et al. 2018; Smail 2024; Wang et al. 2024). These observations have consistently found abundant bright submm/mm sources indicating intense star formation enshrouded by dust in protoclusters. While single-dish surveys provide a wide field-of-view (FOV), which is needed to map protocluster fields, the coarse angular resolution limits the achievable sensitivity due to confusion noise and complicates the identification of counterparts. Interferometers offer angular resolution high enough to overcome the confusion limit of single-dish telescopes, but the small FOV necessitates mosaicking, so interferometric protocluster surveys with a significant area are scarce. Existing interferometric surveys of protoclusters (e.g., Umehata et al. 2017; Oteo et al. 2018) are often too shallow to compare with deep blank field mosaics (e.g., González-López et al. 2020; Fujimoto et al. 2024). As such, the effect of protocluster environment on the abundance of fainter and more representative populations of dusty star-forming galaxies remains poorly explored. Investigating the dust content of other populations in protoclusters, such as Ly blobs (e.g., Umehata et al. 2021) and Lyman break galaxies also requires substantially deeper data.
@@
The Special Selected Area 22 (SSA22) protocluster at was first discovered as a peak in redshift distribution of Lyman break galaxies (Steidel et al. 1998), and then associated with a Mpc scale structure identified by overdensity of Ly emitters (Hayashino et al. 2004; Yamada et al. 2012) distributed along at least three filaments (Matsuda et al. 2005). The large-scale filaments intersect at the core region of the SSA22 protocluster, where numerous massive galaxies (Kubo et al. 2013, 2015, 2016) and AGNs (Lehmer et al. 2009a, b) are seen. Since its discovery, it has been extensively studied from X-ray to radio wavelengths (e.g., Chapman et al. 2004a; Hayashino et al. 2004; Lehmer et al. 2009b; Webb et al. 2009; Kato et al. 2016; Ao et al. 2017). Single-dish submm/mm surveys of the SSA22 field (Geach et al. 2005; Tamura et al. 2009; Umehata et al. 2014) have found substantial star formation activities obscured by dust. Umehata et al. (2017) mapped the central 7 arcmin2 ( Mpc2) area at 1.1 mm using ALMA and revealed an increased number density of bright dusty star-forming galaxies (DSFGs) in the core region. The distribution of member DSFGs and AGNs overlaps with diffuse gas filaments seen in Ly emission (Umehata et al. 2019), suggesting galaxy growth fueled by gas supply from the cosmic web. However, the abundance of fainter DSFGs was not clear because of the small number of detections at an RMS sensitivity of 60 Jy beam-1 (equivalent to a SFR limit of yr-1 at ).
@@ 2.1 ALMA band 6 continuum mosaic
The data were processed by the ALMA pipeline in the CASA (CASA Team et al. 2022) package. Imaging is performed using tclean task with Briggs weighting (robust=0.5), and auto-multithresh masking. We choose the mtmfs deconvolver with scale=[0] because the continuum sources are expected to be at most marginally resolved. The resulting 256.98 GHz (1.17 mm) continuum image is shown in Fig. 1. The synthesized beam has FWHM and position angle (PA)=. The RMS noise level is 25 Jy beam-1 before primary beam correction, of the depth of Jy beam-1 map from Umehata et al. (2017).
@@ 3.1 Source detection
In Fig. 2a, we show the false detection rate as a function of peak S/N. Based on these results, we filter the initial source candidates by requiring peak S/N and create the final source catalogue (Appendix A, see also Fig. 1) from 53 sources, corresponding to a spurious rate of . For each detection, we search for a counterpart in Umehata et al. (2017) main and supplementary catalogues (A1–A32) using a matching radius of beam FWHM (, equivalent to of Gaussian). If no match is found, a new ID starting from A33 is assigned in the order of descending flux. All 18 sources in the main catalogue of Umehata et al. (2017) are detected in the new deeper mosaic, but only 3 out of 14 in the supplementary catalogue with S/N–5.0 (or 2 of 8 sources with S/N) are recovered. Considering the poor confirmation of Umehata et al. (2017) catalogue at S/N, we mark the new sources with as tentative detection.
Source position and flux are measured using the CASA imfit task. The detection completeness is estimated by injecting point sources into random blank positions in the image before primary beam correction, and then running the same source detection script. The artificial source is considered recovered if it is detected at peak S/N within from the injected position. The completeness as a function of source flux is shown in Figure 2b. From the results, we also calculate the ratio between measured and true fluxes (Figure 2c). For sources with peak S/N, we divide their 1.17 mm fluxes from imfit task by the median measured-to-true flux ratio at the corresponding S/N to correct for flux boosting.
@@ 3.2 Counterpart identification and source properties
To search for counterparts at optical and near-infrared wavelengths, we first align the images to the ALMA coordinate frame and subtract the background. A gallery of image stamps at the ALMA source position is provided in Appendix C. The ground-based images have a homogenized point spread function (PSF) FWHM of . Photometry on these images is performed in diameter circular aperture centered on the ALMA position and corrected for aperture loss using the point spread function model in each band. Flux errors are estimated as the standard deviations of fluxes from 1000 randomly placed apertures at blank positions. Fluxes in Spitzer IRAC bands are measured from point source decomposition of the image (Hsieh et al. 2012) in the same aperture as optical images, which allows the source centroid to wander by up to . In Table 3, sources with S/N in any of the , IRAC1 and IRAC2 bands (median depth = 0.3 Jy, 0.3 Jy and 0.4 Jy) are marked as NIR detected. Three sources (A33, A58 and A64) have a nearby NIR source located away from the ALMA position and no photometry. They are marked as potential lensed sources or no counterpart (L/N) in our catalogue.
@@
The effective area as a function of flux is defined as the area in the mosaic where a source can be detected with S/N as shown in Figure 2b. We evaluate the number count from 0.06 to 6.3 mJy in six bins placed evenly in logarithmic space. Field-only number counts are also calculated by assigning a weight of zero to confirmed SSA22 members with spectroscopic redshift . To estimate uncertainty, we resample the flux column of the source table (Appendix A) 1000 times, assuming the flux error follows a Gaussian distribution, and add the resulting standard deviation of the number count in quadrature to the Poisson error as the final uncertainty. The results are listed in Table 1 and plotted in Figure 4, together with a collection of literature results of blank fields (Karim et al. 2013; Ono et al. 2014; Carniani et al. 2015; Hatsukade et al. 2016, 2018; Stach et al. 2018; González-López et al. 2020; Simpson et al. 2020; Béthermin et al. 2020; Gómez-Guijarro et al. 2022; Chen et al. 2023; Fujimoto et al. 2024) and protocluster field (Zhang et al. 2024). The literature fluxes are scaled to 1.17 mm by assuming a modified black body with dust temperature K, emissivity , and median redshift of the DSFG population (e.g., Gómez-Guijarro et al. 2022; Dudzevičiūtė et al. 2020). The resulting number count shows consistency with the field average for flux density mJy but obvious excess at brighter fluxes.
With a similar Monte Carlo method, we also compute the rest-frame 250 m luminosity function for the core region of the SSA22 protocluster (, Matsuda et al. 2005). Assuming a modified black body with dust temperature K, emissivity , and redshift , the 1.17 mm (rest-frame 285 m) luminosity is divided by 0.72 to convert to that at rest-frame 250 m. Protocluster membership is based on spectroscopic redshifts or the probability of photometric redshifts (see the next section) falling into the range of the core region. We resample the catalogue 40,000 times and find the probability that one or more photometric redshift sources are protocluster members to be . However, we note that membership based on photometric redshift can be unreliable: the two sources (A34 and A49) with photometric redshift closest to 3.09 lie in the deep CO(3-2) footprint but are undetected while their continuum fluxes are comparable to A42 at . As such, we mainly consider the results based on spectroscopic membership. The monochromatic luminosity function is listed in Table 2 and plotted in Fig. 4.
@@ 4.2 The rest-frame FIR luminosity function
In Fig. 4, we can see that the rest-frame 250 m luminosity function of the SSA22 protocluster core region is dex above that in coeval blank field (e.g., Lapi et al. 2011; Koprowski et al. 2017; Gruppioni et al. 2020; Fujimoto et al. 2024). Such orders of magnitude higher obscured star formation rate densities than those in blank fields have also been observed in other high-redshift protoclusters (e.g., Oteo et al. 2018; Lacaille et al. 2019; Hill et al. 2020; Zhang et al. 2024). In line with what has been seen in number count, the luminosity function shows evidence of a declining (or at least flat) faint end, suggesting a lack of faint DSFGs in the SSA22 protocluster core region: if the functional form is the same as blank fields, then by scaling the field luminosity function to match the two brightest flux bins of SSA22, we will expect much more sources at S mJy than what has been observed. The overall shape of the rest-frame 250 m luminosity function of the SSA22 protocluster core region is reminiscent of the Gaussian-shape bright end of the optical luminosity function of low-redshift galaxy clusters (e.g., Lobo et al. 1997; Wilson et al. 1997; Bernardi et al. 2003). This points to an evolutionary link between protocluster and local galaxy clusters and supports the idea that the massive early-type galaxies in local clusters have formed the bulk of their stellar mass during star formation events enshrouded by dust at (e.g., Stach et al. 2017; Cooke et al. 2019; Smail 2024). Finally, it is notable that recent high-resolution JWST and ALMA observations of DSFGs in the SSA22 protocluster core have revealed evidence of growing bulge components behind dust obscuration (Umehata et al. 2025), which again suggests that we are witnessing the early formation phase of the massive end of the galaxy population in a protocluster.
@@ 4.3 Stellar mass and SFR of the SSA22 protocluster member galaxies
In Fig. 5, the SSA22 protocluster member galaxies have SFR overlapping with the SFMS, but the distributions of and SFR are different from field DSFGs. The differences are at significance levels from a two-sided KS test except for the distributions of SSA22 protocluster members and GOODS-ALMA DSFGs, which differ at . Overall, SSA22 protocluster galaxies have lower SFR and smaller , and we note that it is unlikely that the differences are driven by the choice of SED fitting codes because their results usually agree within error (Pacifici et al. 2023). Instead, it can be driven by the fainter fluxes of our sample: the median 1.17 mm fluxes are 0.8, 1.3 and 2.0 mJy for SSA22 protocluster member, GOODS-ALMA and AS2UDS+AS2COSMOS DSFGs, respectively. The deeper flux limit of our survey can pick up lower SFR and less massive systems, and the median of each sample (, , ) roughly scales as the median 1.17 mm fluxes. Since the submm/mm fluxes reflects a combination of dust mass and SFR (), this is the expected behavior if we further assume and consider the fact that none of the SSA22 protocluster member galaxies have SFR above SFMS like the field DSFGs. As is a proxy to total gas mass (e.g., Scoville et al. 2016; Hodge & da Cunha 2020), the above results suggest the current star formation in SSA22 protocluster member galaxies occurs in a self-regulated way powered by a large amount of fresh gas supply at the cosmic web node (Umehata et al. 2019), rather than short-lived starbursts induced by major merger. In addition, as will be presented in a future paper, [C ii] observations reveal that the majority of bright SSA22 protocluster member DSFGs are regular rotating disks with widespread star formation (see also, Umehata et al. 2025) instead of late-stage major merger, and the brightest one, A1, is a massive spiral galaxy (Umehata et al. 2024). These observations further support our argument about secular star formation. A similar conclusion has also been drawn for the Spiderweb protocluster (Pérez-Martínez et al. 2024).